OpenStudy (anonymous):

Help! I need an answer for the integral lower limit -1 upper limit 1 x(x^2 +1)^5 I got an answer of 1/12 (x^2+1)^6 +C

OpenStudy (anonymous):

You have the correct antiderivative, but the answer you have isn't the value of the definite integral \(\large\displaystyle \int_{-1}^1 x(x^2+1)^5\,dx\). Using the work you did, we have that \[\large \int_{-1}^1 x(x^2+1)^5\,dx = \left.\left[\frac{1}{12}(x^2+1)^6\right]\right|_{-1}^1\] and now use the fundamental theorem of calculus to evaluate this. Does this make sense? :-) Once you do this, I'll show you an even shorter way to do this problem. :-)

OpenStudy (anonymous):

I'm not sure. I'm so confused right now! Can you show me?

OpenStudy (anonymous):

the shorter way.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You do need to be familiar with the fundamental theorem of calculus anyways, so first note that \[\large \begin{aligned} \left.\left[\frac{1}{12}(x^2+1)^6\right]\right|_{-1}^1 &= \left(\frac{1}{12}(1^2+1)^6\right) - \left(\frac{1}{12}((-1)^2+1)^6\right)\\ &= \frac{2^6}{12} - \frac{2^6}{12}\\ &= 0\end{aligned}\] ---------------- The shorter way is to make an observation about the function you're integrating. Take \(\large f(x) = x(x^2+1)^5\). Note that \[\large f(-x) = -x((-x)^2+1)^5 = -x(x^2+1)^5=-f(x)\] Thus the function we're integrating is ODD. Then, if we evaluate the definite integral of an odd function \(\large f(x)\) over a symmetric interval \(\large [-a,a]\), the result is ZERO, i.e. \[\large \int_{-a}^a f(x)\,dx =0\] Hence, in your problem, you're integrating the odd function \(\large f(x)=x(x^2+1)^5\) over the interval \(\large [-1,1]\). Therefore, you have \[\large \int_{-1}^1x(x^2+1)^5\,dx = 0\] by the property I mentioned above. There is a similar property for even functions: If \(\large f(x)\) is even, then \[\large \int_{-a}^a f(x)\,dx = 2\int_0^a f(x)\,dx.\] I hope this makes sense!